# CAN QUANTUM THEORY BE UNDERPINNED BY A NON-LOCAL HIDDEN VARIABLE THEORY ?

### v1

@@ 3.4 Quantum Entanglement and EPR Steering ?
Bell locality violation also implies that EPR steering and quantum entanglement both occur, however quantum entanglement or EPR steering can also occur for some Bell local states ([13], [9], [14], [15], [16]). Although this paper is focused on tests for Bell locality violation, Bell tests for quantum entanglement or EPR steering are also important ([16]). For completeness, a brief outline of how LHVT Bell states can be divided into three categories with different featues for Quantum Entanglement and EPR Steering is set out in Appendix A. The violation of other Bell inequalities for LHVT involving local hidden quantum states (LHS) [9] show whether or not Quantum entanglement or EPR steering is occuring.
Bell inequalities also exist for multi-partite systems -- GHZ states [17] or for measurements at three different times -- Leggett-Garg inequalities ([18]).  A recent review dealing with Bell correlations in macroscopic and mesoscopic systems is in Ref.[19]
@@ 4 CGLMP BELL INEQUALITIES
The CGLMP Bell inequalities will now be discussed - first for the local hidden variable theory case and then for  the non-local hidden variable theory case.
4.1 LHVT Case
@@
A bipartite system with two observables per sub-system, based on HVT joint probability $P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2})\equiv C(j,k,l,m)$ for all four observables $\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2}$ ([5], [6]) is considered. To shorten the notation we may write $\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}\equiv j,k,l,m$. In this bipartite case there are the same number $d$ of outcomes for each observable listed $10,1,...,d-$ 1 for each observable $\ j,k,l,m$..
@@
The joint probability $P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2})$ applies to the overall result of four different sets of measurements, each involving one observable for each sub-system. These measurement choices are ($\Omega_{A1},\Omega_{B1}$), ($\Omega_{A2},\Omega_{B1}$), ($\Omega_{A2},\Omega_{B2}$) and ($\Omega_{A1},\Omega_{B2}$). The LHVT expression is given by
@@
The probability for joint outcomes of one observable for each sub-system satisfy LHVT conditions [20]. A typical case is $P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1})$ $=\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)P(\alpha_{j}|\Omega_{A1},\lambda)P(\beta_{l}|\Omega_{B1},\lambda)$.
@@
$P(\Omega_{A1}{\small=}\Omega_{B1})+P(\Omega_{B1}{\small=}\overline{\Omega_{A2}{\small+}1})+P(\Omega_{A2}{\small=}\Omega_{B2})+P(\Omega_{B2}{\small=}\Omega_{A1})\leq 3$ (10)
This Bell inequality based on the LHVT expression is violated for maximally entangled state
@@
$\left|\Psi\right\rangle=\sum_{j=0}^{d-1}\left|j\right\rangle_{A}\otimes\left|j\right\rangle_{B}$ (11)
@@
for observables whose eigenstates for eigenvalues $k$, $l$ in the case of observables $\Omega_{Aa}$ , $\Omega_{Bb}$ ($a,b=1,2$) are:
@@
with a similar result for $\left\langle l\;|l^{\ast}\right\rangle_{B,b\text{ .}}$ell non-locality for above CGLMP Bell inequalitity occurs for $d=2$.
@@
where we have used $\sin^{2}(\pi(k-l+\theta_{a}+\phi_{b}))=\{\sin\pi(k-l)\cos\pi(\theta_{a}+\phi_{b})+(\cos\pi(k-l)\sin\pi(\theta_{a}+\phi_{b})\}^{2}=\{0+(\pm 1)(\pm 1/\sqrt{2})\}^{2}$, noting that the only $\theta_{a}+\phi_{b}$ values are $1/4,-1/4,3/4,1/4$ and $\sin\pi(1/4)=-\sin\pi(-1/4)=\sin\pi(3/4)=1/\sqrt{2}$.

### v2

@@ 2.5 Quantum Entanglement and EPR Steering ?
Bell locality violation also implies that EPR steering and quantum entanglement both occur. However quantum entanglement or EPR steering can also occur for some Bell local states ([20], [16], [21], [22], [23]). Although this paper is focused on tests for Bell locality violation, Bell tests for quantum entanglement or EPR steering are also important ([23]). After all, issues regarding these effects were part of the motivation for the search for a theory to underpin quantum theory. For completeness, a brief outline of how LHVT Bell states can be divided into three categories with different featues for Quantum Entanglement and EPR Steering is set out in Appendix A. The violation of other Bell inequalities for LHVT based on this categorisation involving local hidden quantum states (LHS) [16] show whether or not Quantum entanglement or EPR steering is occuring.
Bell inequalities also exist for multi-partite systems such as in GHZ states [24] or for measurements at three different times as in  the  Leggett-Garg inequalities ([25]).  A recent review dealing with Bell correlations in macroscopic and mesoscopic systems is in Ref.[26]
@@ 3 CGLMP BELL INEQUALITIES
The CGLMP Bell inequalities will now be discussed - both for the  local hidden variable theory case and for the non-local hidden variable theory case. Here we consider the case of a bipartite system in which the two sub-systems may be spatialy separated, but now.we consider two observables per sub-system.
3.1 General CGLMP Considerations
@@
A bipartite system with two observables per sub-system, based on a HVT joint probability $P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2})\equiv C(j,k,l,m)$ for all four observables $\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2}$ ([7], [8]) is considered. To shorten the notation we may write $\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}\equiv j,k,l,m$. In this bipartite case there are the same number $d$ of outcomes for each observable listed $10,1,...,d-$ 1 for each observable listed as $\ j,k,l,m$..
CGLMP ([7]) wrote their paper for a deterministic version of HVT. Here the hidden variables are the outcomes $\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}$ themselves. However,they pointed out that their theory also has a probabilistic version,- in which hidden variables $\lambda$ only determine the probabilities $P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$ of the outcomes. for particular $\lambda$.  In this probabilistic case we would have
$P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2})$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$
For each HV choice $\lambda$ the probabilities for all outcomes must add to unity, so $\sum_{jklm}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)=1$. This form for $P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2})$ applies for both the LHVT and NLHVT situations, with the (below) LHVT form (17) involving a factorisation of $P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$.
There is no reason why the HVT cannot be a NLHVT whose basic probability is given by Eq (LABEL:Eq.NLHVTCGLMP) and which involves two observables per sub-system. After all, this is consistent with the idea of realism in which the hidden variables are the actual outcomes $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}$ of the observables $\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2}$ in the deternministic version or their probabilities $P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$ in the probabilistic version with  hidden variables $\lambda$. Such an approach is the basis of the.CGLMP paper. In both probabilistic and deterministic versions the outcomes have a real existence prior to any measurements, as Einstein wanted in any theory underpinning quantum theory. Also in hidden variable theory, the observables are not non-commuting Hermitian operators, so it is legitimate to have an approach where all obervables (commuting or non -commuting in quantum theory) have simultaneous outcomes. In Bohmian Mechanics the positions and momenta of every particle is ascribed a real existence. Of course for comparison with quantum theory or experiment we must construct expressions that correspond to quantitities obtainable from quantum theory formulae or which can  be measured. However this can be achieved - as for quantities such as $\left\langle\Omega_{Ai}\otimes\Omega_{Bj}\right\rangle$, which in quantum theory are given by $Tr(\widehat{\Omega}_{Ai}\otimes\widehat{\Omega}_{Bj}\;\widehat{\rho})$ or in experiment by repeated measurements on each pair of observables $\Omega_{Ai}$ and $\Omega_{Bj}$.to determine the mean value of the product of the outcomes.
CGMP also only considered cases where the outcomes for one of the observables for each sub-system was considered - $\Omega_{A1}$ or $\Omega_{A2}$ for sub-system $A$ and $\Omega_{B1}$ or $\Omega_{B2}$ for sub-system $B.$ We restrict ourselves to this situation, and hence the only HVT marginal probabilities we consider (in the general probabilistic version) are
$P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1})$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\mathop{\displaystyle\sum}\limits_{\alpha_{k}\beta_{m}}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$
$P(\alpha_{k},\beta_{l1}|\Omega_{A2},\Omega_{B1})$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\mathop{\displaystyle\sum}\limits_{\alpha_{j}\beta_{m}}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$
$P(\alpha_{j},\beta_{m}|\Omega_{A2},\Omega_{B2})$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\mathop{\displaystyle\sum}\limits_{\alpha_{k}\beta_{l}}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$
$P(\alpha_{k},\beta_{m}|\Omega_{A1},\Omega_{B2})$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\mathop{\displaystyle\sum}\limits_{\alpha_{j}\beta_{l}}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$
So far we have not specified whether the CGLMP  Bell inequalities we consider are based on a local or a non-local version of HVT. [7] state that they are considering a local version of HVT. This statement is rather puzzling as CGLMP do not express $P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$ or $C(j,k,l,m)$ as the product of separate sub-system probabilities - which could be either as $P(\alpha_{1},\alpha_{2}|\Omega_{A1},\Omega_{A2};\lambda)P(\beta_{1},\beta_{2}|\Omega_{B1},\Omega_{B2};\lambda)$ or as $P(\alpha_{j}|\Omega_{A1},\lambda)P(\alpha_{k}|\Omega_{A2},\lambda)P(\beta_{l}|\Omega_{B1},\lambda)P(\beta_{m}|\Omega_{B2},\lambda)$. This issue is discussed in an earlier paper on the CGLMP inequalities [27]. It was pointed out there that a theorem by Fine [28] shows that the marginal probabilities (such as $P(\alpha_{1},\beta_{1}|\Omega_{A1},\Omega_{B1})$ in Eq (LABEL:Eq.MargProb2) can be written in LHVT form. As these marginal probabilities are used to evaluate the quantity $S$ (see Eq (30) ) in the CGLMP Bell inequality, it could be claimed that CGLMP is based on  LHVT and hence a violation of the inequality shows quantum theory can not be underpinned by LHVT. That may be the case, but what we now show is that the CGLMP Bell inequality can also be established using NLHVT (just as for CHSH) - and hence the inequality - being inconsistent with quantum theory - shows that quantum theory is also not underpinned by this non-local hidden variable theory.
@@
The joint probability $P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2})$ in the CGLMP Bell inequality (30) is applied in determining the overall result for four different sets of measurements, each involving one observable for each sub-system. These measurement choices are ($\Omega_{A1},\Omega_{B1}$), ($\Omega_{A2},\Omega_{B1}$), ($\Omega_{A2},\Omega_{B2}$) and ($\Omega_{A1},\Omega_{B2}$).
For completeness - using the locality condition (LABEL:Eq.NLHVTCGLMP)for $P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$ , we see that the LHVT expression for the joint probability is given by
@@
The marginal probabilities for joint outcomes of one observable for each sub-system satisfy LHVT conditions [28]. A typical marginal probability for LHVT is $P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1})$ $=\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)P(\alpha_{j}|\Omega_{A1},\lambda)P(\beta_{l}|\Omega_{B1},\lambda)$ which is obviously of LHVT form. Note that $\mathop{\displaystyle\sum}\limits_{\alpha_{k}\beta_{m}}P(\alpha_{k}|\Omega_{A2},\lambda)P(\beta_{m}|\Omega_{B2},\lambda)=1$. For the NLHVT the results for the marginal probabilites are just given by (LABEL:Eq.MargProb2).
@@
A typical CGLMP Bell inequality is
$S=P(\Omega_{A1}{\small=}\Omega_{B1})+P(\Omega_{B1}{\small=}\overline{\Omega_{A2}{\small+}1})+P(\Omega_{A2}{\small=}\Omega_{B2})+P(\Omega_{B2}{\small=}\Omega_{A1})\leq 3$ (18)
3.2 NLHVT Proof of CHSH Bell Inequality
The proof of the CHSH Bell inequality based on local hidden variable theory is well known [6], and is set out in Ref  [9]. However, as the CHSH Bell inequality involves two observables per sub-system $\Omega_{A1},\Omega_{A2}$  for sub-system $A$ with outcomes $\alpha_{1},\alpha_{2}$ and similarly for sub-system $B,$ it is of interest to consider the situation where a general non-local hidden variable theory applies. Here there is a non-local hidden variable theory joint probability $P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2})$ for the measurement of all four observables - even though each pair of sub-system observable cannot be simultaneously measured according to quantum theory. Probability conservation gives $\sum_{\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}}P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2})$ $=1$. The question is: Does the CHSH Bell inequality still apply ?.
In this case for a probabilistic HVT
$P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2})$ (19)
$=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$ (20)
The marginal probabilities for measurement outcomes for one observable for each sub-system would be
$P(\alpha_{1},\beta_{1}|\Omega_{A1},\Omega_{B1})$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\mathop{\displaystyle\sum}\limits_{\alpha_{2}\beta_{2}}P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$
$P(\alpha_{2},\beta_{1}|\Omega_{A2},\Omega_{B1})$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\mathop{\displaystyle\sum}\limits_{\alpha_{1}\beta_{2}}P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$
$P(\alpha_{1},\beta_{2}|\Omega_{A2},\Omega_{B2})$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\mathop{\displaystyle\sum}\limits_{\alpha_{2}\beta_{1}}P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$
$P(\alpha_{2},\beta_{2}|\Omega_{A1},\Omega_{B2})$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\mathop{\displaystyle\sum}\limits_{\alpha_{1}\beta_{1}}P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)$
where all four choices of the pairs of sub-system observables $(\Omega_{Aa},\Omega_{Bb})$ are considered. Note the sums over the unrecorded observable outcomes. These expressions are as (LABEL:Eq.MargProb2) with a notation change.
Typical terms in the expression for $S$ can then be obtained. For the first term
@@ 3.2 NLHVT Proof of CHSH Bell Inequality
$\left\langle\Omega_{A1}\otimes\Omega_{B1}\right\rangle_{LHVT}$ $=$ $\sum_{\alpha_{1},\beta_{1},}\{\alpha_{1}\beta_{1}\}P(\alpha_{1},\beta_{1}|\Omega_{A1},\Omega_{B1})$
$=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\sum_{\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}}P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\{\alpha_{1}\beta_{1}\}$
The other terms are derived similarly.
The quantity $S$ in the CHSH inequality would then be given by
$S$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\sum_{\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}}P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\{\alpha_{1}\beta_{1}+\alpha_{1}\beta_{2}+\alpha_{2}\beta_{1}-\alpha_{2}\beta_{2}\}$
$=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\sum_{\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}}P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\{\alpha_{1}(\beta_{1}+\beta_{2})+\alpha_{2}(\beta_{1}-\beta_{2})\}$
In the CHSH case where $\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2}$ are all components of Pauli spin observables in spin $1/2$ sub-systems, all the outcomes are either $+1$ or $-1$. In which case $(\beta_{1}+\beta_{2})$ can either be $-2$, 00, 00 or $+2$ and for these situations $(\beta_{1}-\beta_{2})$ will be 00, $+2$, $-2$ or 00 respectively. Thus one of the two factors $(\beta_{1}+\beta_{2})$ or $(\beta_{1}-\beta_{2})$ will be zero whilst the other has a magnitude of $2$ The magnitude of both $\alpha_{1}$ and $\alpha_{2}$ is $1$ Thus in all cases |$|\{\alpha_{1}(\beta_{1}+\beta_{2})+\alpha_{2}(\beta_{1}-\beta_{2})\}|\;=2$.
Hence
$|S|\,$$\leq$ $\,\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\sum_{\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}}P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\times 2$ (24)
$\leq$  $\,2\text{. }$
as required - using the magnitude of a sum being less than or equal to the sum of the magnitudes.
The CHSH Bell inequality can also be proved based on the LHVT form (17) for the joint probability with two observables per sub-system, where $P(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)=P(\alpha_{1}|\Omega_{A1},\lambda)P(\alpha_{2}|\Omega_{A2},\lambda)P(\beta_{1}|\Omega_{B1},\lambda)P(\beta_{2}|\Omega_{B2},\lambda).$ So it is only for the one observable per sub-system form (4) for the NLHVT where the CHSH Bell inequality proof fails.
Thus the CHSH inequality also holds for this type of NLHVT, so its violation for the singlet state listed in Section 2.3 shows that this type of non-locality is not consistent with quantum theory. Hence we find the new result that quantum theory cannot be underpinned by this type of NLHVT. Evidently, the violation of the CHSH Bell inequality also rules out a very general type of NLHVT as well ! We will find that the same outcome applies when the CGLMP Bell inequalities are considered.
3.3 Observables and Quantum State
We now introduce the quantum theory observables that will be used in the CGLMP inequalities, along with the quantum state which will be used for comparison with HVT predictions - both local and non-local. .
The Bell inequality (18) based on the LHVT expression is violated for the quantum theory maximally entangled state
$\left|\Psi\right\rangle=\frac{1}{\sqrt{d}}\sum_{j=0}^{d-1}\left|j\right\rangle_{A}\otimes\left|j\right\rangle_{B}$ (25)
@@ 3.3 Observables and Quantum State
for quantum theory observables whose eigenstates for eigenvalues $k$, $l$ in the case of observables $\Omega_{Aa}$ , $\Omega_{Bb}$ ($a,b=1,2$) are:
@@
with a similar result for $\left\langle l\;|l^{\ast}\right\rangle_{B,b\text{ .}}$ell non-locality for above CGLMP Bell inequalitity (18occurs for $d=2$.
@@
where we have used $\sin^{2}(\pi(k-l+\theta_{a}+\phi_{b}))=\{\sin\pi(k-l)\cos\pi(\theta_{a}+\phi_{b})+(\cos\pi(k-l)\sin\pi(\theta_{a}+\phi_{b})\}^{2}=\{0+(\pm 1)(\pm 1/\sqrt{2})\}^{2}$, noting that the only $\theta_{a}+\phi_{b}$ values are $1/4,-1/4,3/4,1/4$ and $\sin\pi(1/4)=-\sin\pi(-1/4)=\sin\pi(3/4)=1/\sqrt{2}$.  Also, $(k-l)$ is always an integer.